# Fast post-process Bayesian inference with Variational Sparse Bayesian Quadrature

### v2

@@ 4.5 Additional analyses
We ran again all our experiments using optimization traces from a different optimization algorithm based on a hybrid Bayesian optimization approach (Garnett, 2023), namely Bayesian Adaptive Direct Search (bads; Acerbi and Ma, 2017). bads is a state-of-the-art optimization algorithm with wide application in computational neuroscience and other fields. We found the performance of vsbq with bads is almost identical to cma-es for three out of four benchmark problems (Two Moons, multivariate Rosenbrock-Gaussian, and Bayesian timing model), and slightly worse for the multisensory causal inference problem. See Supplementary Material C.4 for the full results and their discussion.
@@
Ideally, running a post-process inference method should only take a relatively short time (e.g., a few minutes), so we performed a detailed comparison of the runtimes of different algorithms on different computational architectures (CPU and GPU). Overall, we found that vsbq takes several minutes on CPU and 1-3 minutes on GPU across the various problems we considered, meeting the desiderata of a post-process technique. nnr can take considerably longer, mainly due to multiple training runs required for hyperparameter selection. Full results are reported in Supplementary Material C.6.
@@
In this paper, we presented results obtained by approximating the target posterior density from the surrogate of the log density (Figure 1) via variational inference. A natural alternative is to instead run mcmc on the surrogate log-density (Rasmussen, 2003; Nemeth and Sherlock, 2018; Järvenpää et al., 2021). In Supplementary Material C.7, we provide experimental results showing that mcmc can underperform in this setting and discuss the reasons.
@@ 5 Discussion
In this paper, we introduced the framework of post-process, black-box Bayesian inference, and proposed a specific post-process algorithm, Variational Sparse Bayesian Quadrature (vsbq). By recycling evaluations from previous MAP optimization runs, vsbq enables full Bayesian inference at a limited additional cost. Our method produces high-quality solutions as demonstrated on a series of benchmarks. While our results show the promise of post-process Bayesian inference, and of vsbq in particular, several aspects remain open for future research. We discuss the limitations of our approach and potential future work below.
@@ 5.1 Limitations and future work
Arguably, post-process inference is constrained to low-dimensional problems for several reasons: as the dimension increases, pre-existing evaluations from MAP optimizations become less likely to cover the majority of the posterior mass (Vershynin, 2018, Chapter 3.3.3). While a larger number of evaluations might expand the range of tractable dimensions, the curse of dimensionality poses challenges in approximating a function without additional structural assumptions. A potential solution to address the issue of lack of coverage is active sampling, that is the ability to acquire new log-density evaluations where needed (Osborne et al., 2012; Acerbi, 2020; Järvenpää et al., 2021; De Souza et al., 2022). Active sampling would need one or multiple rounds of ad-hoc function evaluations and subsequent surrogate model updates, which would substantially increase the post-processing time (De Souza et al., 2022). Finding efficient methods for active learning for post-process inference is left for future research. Despite the above limitations, the relatively cheap cost of our method makes it still valuable for quickly constructing a tractable initial approximation of the posterior, potentially useful to inform subsequent runs of mcmc or other inference methods (Zhang et al., 2022).
@@
An important feature of approximate inference methods is the availability of diagnostics to assess the reliability of inference (Vehtari et al., 2021; Yao et al., 2018). A general-purpose inference diagnostic consists of posterior-predictive checks, i.e., testing that synthetic data generated from parameters sampled from the posterior are compatible with the actual data (Gelman et al., 2013). As an additional diagnostic, vsbq provides the standard deviation of the elbo, $\textsc{elbo}_{\text{sd}}$, which can be calculated via sparse Bayesian quadrature (Eq. 23). $\textsc{elbo}_{\text{sd}}$ reflects the uncertainty of the sparse GP prediction in regions where the variational posterior has non-negligible mass. Therefore, it can serve as a useful diagnostic in that solutions with $\textsc{elbo}_{\text{sd}}\gg 1$ should not be trusted. However, as is often the case with inference diagnostics, a small $\textsc{elbo}_{\text{sd}}$ does not guarantee the validity of the approximate variational posterior (Acerbi, 2018). As a practical validation approach, we recommend running vsbq multiple times by dropping for example 20% of the training set and checking the consistency of the approximate posteriors via a form of cross-validation. Visualization of the posterior together with the locations of the evaluated points is also helpful for validating that the approximate posterior is supported by actual evaluations, as opposed to escaping from the training points region, leading to ‘hallucinated’ posterior regions (De Souza et al., 2022). We further discuss this ‘hallucination’ problem in Supplementary Material C.7. Finally, if additional exact log-density evaluations are possible, one can leverage Pareto smoothed importance sampling (Vehtari et al., 2015) for correcting and validating the approximate posterior (Yao et al., 2018).
@@ Acknowledgments
Martin Jørgensen and Luigi Acerbi were supported by the Research Council of Finland (grants number 356498 and 358980 to Luigi Acerbi).
@@ Appendix C Experiment details and additional results
C.1 Noise shaping and choice of hyperparameters
@@
Following these principles, we propose form used in the main text (Eq. 21),
@@
The number of variational components $K$ in both vsbq and nnr is 50, which is also the default maximum number of components in vbmc [Acerbi, 2018].
@@ C.2 Further details of procedure and metrics
C.3 Neural network regression
@@
After neural network fitting, we run stochastic variational inference (svi; Kucukelbir et al., 2017) to compute a tractable posterior approximation from the neural network surrogate. The variational distribution is the same as that used for vsbq, i.e., a mixture of $K=50$ multivariate normal distributions. Unlike vsbq where we can compute the expected log joint in the elbo analytically, with a neural network surrogate, we have to estimate the value and gradient of the expected log joint. We use reparametrization tricks [Kingma and Welling, 2013] to get an unbiased estimate for the gradient of the expected log joint value. Apart from the difference in computing the expected log joint, all the other steps (optimization iterations, variational distribution initialization, etc.) stay the same as the variational inference part in vsbq.
C.4 MAP estimates via Bayesian Adaptive Direct Search (BADS)
@@
We provide the results for vsbq with cma-es and bads in Table S1, S2 and Figure S2, S3. From the tables and figures, we can see that the performance of vsbq with bads and cma-es is almost identical for three out of four benchmark problems (Two Moons, multivariate Rosenbrock-Gaussian, and Bayesian timing model). Instead, we found that vsbq performs less effectively with traces from bads in the multisensory causal inference model, compared to cma-es. Still, vsbq (bads) is comparable to the noiseless Laplace approximation (laplace) even in the presence of large amounts of log-density evaluation noise.
@@
Figure S2: Performance on Bayesian timing model with traces from different optimization algorithms. Median $\Delta$LML loss (left), MMTV (middle), and GsKL (right) as a function of the log-likelihood noise $\sigma_{\text{obs}}$ for the Bayesian Timing model. Shaded areas are 95% CI of the median and black dashed horizontal lines are the rule-of-thumb thresholds for good performance ($\Delta$LML=1, MMTV =0.2, GsKL =$1/8$). The performance of vsbq is virtually identical, regardless of the source of evaluations, whether cma-es or bads.
Figure S3: Performance on multisensory causal inference model with traces from different optimization algorithms. Median $\Delta$LML loss (left), MMTV (middle), and GsKL (right) as a function of the log-likelihood noise $\sigma_{\text{obs}}$ for the Bayesian Timing model. Shaded areas are 95% CI of the median and black dashed horizontal lines are the rule-of-thumb thresholds for good performance ($\Delta$LML=1, MMTV =0.2, GsKL =$1/8$). vsbq with traces from bads performs worse than vsbq with cma-es, but still comparable to the (noiseless) laplace.
C.5 Visualization of posteriors
We visualize posterior distributions as ‘corner plots’, i.e., a plot with 1D and all pairwise 2D marginals. For visualization of individual posteriors obtained by the algorithms, for all problems, we report example solutions obtained from a run with the same random seed (see Figure S4, S6, S8, and S10).
@@ C.5 Visualization of posteriors
Figure S4: Two Moons bimodal posterior visualization. The orange density contours and points in the sub-figures represent the posterior samples from vsbq, nnr, and laplace, while the black contours and points denote ground truth samples. vsbq and nnr perfectly capture the ground-truth bimodal posterior, while laplace is limited to one mode.
@@
Figure S6: Multivariate Rosenbrock-Gaussian posterior visualization. The orange density contours and points in the sub-figures represent the posterior samples from vsbq, nnr, and laplace, while the black contours and points denote ground truth samples. Both vsbq and nnr capture very well the complex shape of the distribution, while in this example laplace fails.
@@
Figure S8: Bayesian timing model posterior visualization. $\sigma_{\textrm{obs}}=3$ for vsbq and nnr, $\sigma_{\textrm{obs}}=0$ for laplace. The orange density contours and points in the sub-figures represent the posterior samples from vsbq, nnr, and laplace, while the black contours and points denote ground truth samples. Both vsbq and nnr capture the shape of the posterior in the presence of observation noise. laplace obtains a reasonable approximation, for the noiseless case.
@@
Figure S10: Multisensory causal inference model posterior visualization. $\sigma_{\textrm{obs}}=3$ for vsbq and nnr, $\sigma_{\textrm{obs}}=0$ for laplace. The orange density contours and points in the sub-figures represent the posterior samples from vsbq, nnr, and laplace, while the black contours and points denote ground truth samples. $(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})=(\sigma_{\text{vis}}(c_{\text{low}}),\sigma_{\text{vis}}(c_{\text{med}}),\sigma_{\text{vis}}(c_{\text{high}}),\sigma_{\text{vest}},\lambda,\kappa)$. Both vsbq and nnr obtain a reasonable approximation of the complex posterior under noisy evaluations, with nnr yielding a more faithful approximation for this random seed. laplace fails to capture the posterior well, despite it being noiseless.
C.6 Runtime analysis
In this section, we present an analysis of the runtime for our method and baselines. Such an analysis is important for a post-process inference method, which ideally should only take a relatively short time (e.g., a few minutes). For vsbq and nnr, we report the wall-clock runtime based on 5 independent runs with different training data sets (i.e., traces from the MAP optimizations). For laplace, the runtime is based on 5 runs for numerically estimating the Hessian matrix. We measure the runtime on both CPU and GPU platforms. [^3] It is worth noting that, in the case of GPU runtime measurement, only the sparse GP and neural network fitting were carried out with a GPU. Due to our current implementation, the stochastic variational inference optimization part for vsbq and nnr was always conducted using the CPU. Future efforts to port the stochastic variational inference procedure completely to GPU could further enhance the computational performance of both methods. Specifically, we utilize an AMD EPYC 7452 32-core Processor for CPU computations and an NVIDIA A100 for GPU computations.

### v3

@@ 4.5 Additional analyses
We ran again all our experiments using optimization traces from a different optimization algorithm based on a hybrid Bayesian optimization approach (Garnett, 2023), namely Bayesian Adaptive Direct Search (bads; Acerbi and Ma, 2017). bads is a state-of-the-art optimization algorithm with wide application in computational neuroscience and other fields. We found the performance of vsbq with bads is almost identical to cma-es for three out of four benchmark problems (Two Moons, multivariate Rosenbrock-Gaussian, and Bayesian timing model), and slightly worse for the multisensory causal inference problem. See Supplementary Material C.5 for the full results and their discussion.
@@
Ideally, running a post-process inference method should only take a relatively short time (e.g., a few minutes), so we performed a detailed comparison of the runtimes of different algorithms on different computational architectures (CPU and GPU). Overall, we found that vsbq takes several minutes on CPU and 1-3 minutes on GPU across the various problems we considered, meeting the speed desiderata of a post-process technique. nnr can take considerably longer, mainly due to multiple training runs required for hyperparameter selection. bbvi methods are not intended as post-processing approaches, making their runtime less directly relevant. The runtime of bbvi and Laplace approximation methods depends on the cost and the number of target density evaluations. Full results are reported in Supplementary Material C.7.
@@
In the earlier paragraphs, we presented results obtained by approximating the target posterior density from the surrogate of the log density (Figure 1) via variational inference. A natural alternative is to instead run mcmc on the surrogate log-density (Rasmussen, 2003; Nemeth and Sherlock, 2018; Järvenpää et al., 2021; Gammal et al., 2022). In Supplementary Material C.8, we provide experimental results showing that mcmc can underperform in this setting and discuss the reasons.
@@ 5 Discussion
In this paper, we introduced the framework of post-process, black-box Bayesian inference, and proposed a specific post-process algorithm, Variational Sparse Bayesian Quadrature (vsbq). By recycling evaluations from previous MAP optimization runs, vsbq enables full Bayesian inference at a limited additional cost. In this section, we first discuss why surrogate-based approaches like vsbq can be more effective than bbvi, followed by an exploration of the limitations of our method and potential directions for future work.
5.1 Why are surrogate-based approaches more effective than direct variational inference?
As demonstrated across a series of benchmarks, in our black-box setting surrogate-based methods like vsbq and nnr produce high-quality solutions, whereas bbvi– which performs black-box variational inference directly on the target – struggles to effectively fit the target posterior, even when given 10×10$\times$10 × target density evaluations. This disparity arises primarily from the high variance associated with the score function estimator in bbvi. While gradient-based variational inference using the reparameterization trick (such as automatic differentiation variational inference, or advi) empirically reduces variance and is more widely adopted (Titsias and Lázaro-Gredilla, 2014; Kucukelbir et al., 2017; Xu et al., 2019), it requires the target density to be differentiable with respect to the parameters. The black-box nature of the target model makes advi inapplicable in our considered scenarios. Surrogate-based methods address these challenges effectively. Fitting a surrogate model (e.g., a sparse Gaussian process for vsbq or a neural network for nnr) to the target density offers two key advantages:
  1. a.
    These methods can fully leverage all existing target density evaluations from various sources, effectively interpolating and smoothing across (noisy) observations.
  2. b.
    The surrogate model resolves the non-differentiability issue, allowing gradient-based variational inference to be applied to the surrogate density, which is differentiable (e.g., via automatic differentiation) by construction.
@@ 5.2 Limitations and future work
Arguably, post-process inference via vsbq is constrained to low-dimensional problems for several reasons: as the dimension increases, pre-existing evaluations from MAP optimizations become less likely to cover the majority of the posterior mass (Vershynin, 2018, Chapter 3.3.3). Moreover, the MAP estimate may become not meaningful or even ill-defined in certain settings, e.g., when the likelihood is unbounded (Gelman et al., 2013). While a larger number of evaluations might expand the range of tractable dimensions, the curse of dimensionality poses challenges in approximating a function without additional structural assumptions. A potential solution to address the issue of lack of coverage is active sampling, that is the ability to acquire new log-density evaluations where needed (Osborne et al., 2012; Acerbi, 2020; Järvenpää et al., 2021; De Souza et al., 2022). Active sampling would need one or multiple rounds of ad-hoc function evaluations and subsequent surrogate model updates, which would substantially increase the post-processing time (De Souza et al., 2022). Finding efficient methods for active learning for post-process inference is left for future research. Despite the above limitations, the relatively cheap cost of our method makes it still valuable for quickly constructing a tractable initial approximation of the posterior, potentially useful to inform subsequent runs of mcmc or other inference methods (Zhang et al., 2022).
@@
An important feature of approximate inference methods is the availability of diagnostics to assess the reliability of inference (Vehtari et al., 2021; Yao et al., 2018). A general-purpose inference diagnostic consists of posterior-predictive checks, i.e., testing that synthetic data generated from parameters sampled from the posterior are compatible with the actual data (Gelman et al., 2013). As an additional diagnostic, vsbq provides the standard deviation of the elbo, $\textsc{elbo}_{\text{sd}}$, which can be calculated via sparse Bayesian quadrature (Eq. 23). $\textsc{elbo}_{\text{sd}}$ reflects the uncertainty of the sparse GP prediction in regions where the variational posterior has non-negligible mass. Therefore, it can serve as a useful diagnostic in that solutions with $\textsc{elbo}_{\text{sd}}\gg 1$ should not be trusted. However, as is often the case with inference diagnostics, a small $\textsc{elbo}_{\text{sd}}$ does not guarantee the validity of the approximate variational posterior (Acerbi, 2018). As a practical validation approach, we recommend running vsbq multiple times by dropping for example 20% of the training set and checking the consistency of the approximate posteriors via a form of cross-validation. Visualization of the posterior together with the locations of the evaluated points is also helpful for validating that the approximate posterior is supported by actual evaluations, as opposed to escaping from the training points region, leading to ‘hallucinated’ posterior regions (De Souza et al., 2022). We further discuss this ‘hallucination’ problem in Supplementary Material C.8. Finally, if additional exact log-density evaluations are possible, one can leverage Pareto smoothed importance sampling (Vehtari et al., 2015) for correcting and validating the approximate posterior (Yao et al., 2018).
@@ Acknowledgments
Chengkun Li, Martin Jørgensen and Luigi Acerbi were supported by the Research Council of Finland (grants number 356498 and 358980 to Luigi Acerbi).
@@ Appendix C Experiment details and additional results
C.1 Implementation details
In this section, we describe the implementation details of variational sparse Bayesian quadrature (vsbq), including noise shaping, choice of hyperparameters, and variational inference details.
@@ C.1 Implementation details
Following these principles, we propose the form used in the main text (Eq. 21),
@@
Stochastic variational inference.
The number of variational components $K$ in vsbq is 50, which is also the default maximum number of components in vbmc [Acerbi, 2018]. For initialization, we first perform K-means clustering on a subset of training points and initialize each component location $\bm{\mu}_{k}^{(i)}$ around the cluster centers, adding Gaussian noise with a standard deviation $10^{-6}$, for $1\leq k\leq K$. The initial scale of the component $\sigma_{k}\bm{\lambda}^{(i)}$ is set to $10^{-3}$, for each dimension $i$, where $1\leq i\leq D$. For the two moons bimodal problem, we set the number of clusters to 50, using the top $80\%$ high-density points for K-means clustering. For all other problems, the number of clusters is set to 1, with K-means applied to the top $1\%$ of high-density points – effectively computing the mean of the selected points.
As described in the main text, after fitting the sparse GP surrogate, variational inference is conducted by optimizing the surrogate elbo with Bayesian quadrature. In addition, we impose a soft penalty loss during the elbo optimization for bounding the variational parameters (means and scales of the mixture components), as done in Acerbi [2020], to help constrain the variational distribution in the local trust region of the surrogate. The lower and upper bounds are computed based on the training points $\mathbf{X}$. For each dimension $i$, let $\mathbf{X}^{(i)}_{\text{min}}$ and $\mathbf{X}^{(i)}_{\text{max}}$ denote the minimum and maximum value of this dimension. The lower bounds and upper bounds for the $k^{\text{th}}$ component’s mean $\bm{\mu}_{k}^{(i)}$ and log scale $\log\sigma_{k}\bm{\lambda}^{(i)}$ are provided in Table S1.
Parameter Description Lower bound Upper bound
$\bm{\mu}_{k}^{(i)}$ mixture component mean $\mathbf{X}^{(i)}_{\text{min}}$ $\mathbf{X}^{(i)}_{\text{max}}$
$\log\sigma_{k}\bm{\lambda}^{(i)}$ mixture component scale $\log\left[10^{-6}\left(\mathbf{X}^{(i)}_{\text{max}}-\mathbf{X}^{(i)}_{\text{min}}\right)\right]$ $\log\left(\mathbf{X}^{(i)}_{\text{max}}-\mathbf{X}^{(i)}_{\text{min}}\right)$
Table S1: The soft bounds for variation posterior parameters.
For both the mean and scale, the soft penalty loss can be written as,
$\mathbf{1}_{\theta_{k}^{(i)}\leq\text{LB}(\theta_{k}^{(i)})\text{ or }\theta_{k}^{(i)}\geq\text{UB}(\theta_{k}^{(i)})}\cdot\frac{1}{2}\left[\frac{\max\left(\theta_{k}^{(i)}-\text{LB}(\theta_{k}^{(i)}),\text{UB}(\theta_{k}^{(i)})-\theta_{k}^{(i)}\right)}{\tau\left(\text{UB}(\theta_{k}^{(i)})-\text{LB}(\theta_{k}^{(i)})\right)}\right]^{2},$ (S4)
where $\theta_{k}^{(i)}$ represents either the mean $\bm{\mu}_{k}^{(i)}$ or the log scale $\log\sigma_{k}\bm{\lambda}^{(i)}$, and $\tau=0.01$. $\text{LB}(\theta_{k}^{(i)})$ and $\text{UB}(\theta_{k}^{(i)})$ denote the lower and upper bounds, respectively.
@@ C.2 Further details of procedure and metrics
C.3 Black-box variational inference
For black-box variational inference, the score function estimator for the elbo gradient can be written as:
$\begin{split}\nabla_{\bm{\phi}}\mathbb{E}_{\bm{\phi}}\left[\log\frac{p(\mathcal{D}|\mathbf{x})p(\mathbf{x})}{q_{\bm{\phi}}(\mathbf{x})}\right]&=\mathbb{E}_{\bm{\phi}}\left[\nabla_{\bm{\phi}}\log q_{\bm{\phi}}(\mathbf{x})\left(\log p(\mathcal{D}|\mathbf{x})p(\mathbf{x})-\log q_{\bm{\phi}}(\mathbf{x})\right)\right]\\ &=\mathbb{E}_{\bm{\phi}}\left[\nabla_{\bm{\phi}}\log q_{\bm{\phi}}(\mathbf{x})h_{\bm{\phi}}(\mathbf{x})\right],\end{split}$ (S5)
where $h_{\bm{\phi}}(\mathbf{x})=\log p(\mathcal{D}|\mathbf{x})p(\mathbf{x})-\log q_{\bm{\phi}}(\mathbf{x})$. As shown by Ranganath et al. [2014], there are two possible ways to reduce the variance of the score function estimator: the Rao-Blackwellization technique and control variates. The Rao-Blackwellization technique cannot be used since we only have access to the target density while the dependency structure of the black-box target model is assumed unknown. Conversely, we can use control variates. In particular, noting that for any constant $b$, $\mathbb{E}_{\bm{\phi}}[b\nabla_{\bm{\phi}}\log q_{\bm{\phi}}]=0$, $b\nabla_{\bm{\phi}}\log q_{\bm{\phi}}$ can be used as the control variate for reducing the stochastic gradient variance, i.e.,
$\mathbb{E}_{\bm{\phi}}\left[\nabla_{\bm{\phi}}\log q_{\bm{\phi}}(\mathbf{x})h_{\bm{\phi}}(\mathbf{x})\right]=\mathbb{E}_{\bm{\phi}}\left[\nabla_{\bm{\phi}}\log q_{\bm{\phi}}(\mathbf{x})(h_{\bm{\phi}}(\mathbf{x})-b)\right].$ (S6)
An optimal choice of $b$ would require estimating the covariance between $\nabla_{\bm{\phi}}\log q_{\bm{\phi}}(\mathbf{x})h_{\bm{\phi}}(x)$ and $\nabla_{\bm{\phi}}\log q_{\bm{\phi}}$, and the variance of $\nabla_{\bm{\phi}}\log q_{\bm{\phi}}$, which would require extra evaluations on the target density function. To simplify the implementation and comparison to other methods, we instead take an exponential moving average of $h_{\bm{\phi}}(\mathbf{x})$ as the value of $b$, as suggested in the probabilistic programming framework Pyro [Bingham et al., 2018]. The smoothing factor for the exponential moving average is 0.9.
As stated in the main text, we allocate $10\times 3000D$ target density evaluation budget for BBVI. Denoting the number of Monte Carlo samples for gradient estimation with $M_{g}\in\{1,10,100\}$, the total number of optimization iterations is then set to $\frac{10\times 3000D}{M_{g}}$. For all variational distribution families—a Gaussian with diagonal covariance matrix, a Gaussian with full-rank covariance matrix, and a mixture of Gaussians with $K=5$ and $K=50$ components—the Gaussian distribution mean is initialized near the origin, by adding Gaussian noise with a standard deviation $10^{-6}$ and the scales are set to $10^{-3}$.
C.4 Neural network regression
@@
After neural network fitting, we run stochastic variational inference via automatic differentiation (advi; Kucukelbir et al., 2017) to compute a tractable posterior approximation from the neural network surrogate. The variational distribution is the same as that used for vsbq, i.e., a mixture of $K=50$ multivariate normal distributions. Unlike vsbq where we can compute the expected log joint in the elbo analytically, with a neural network surrogate, we have to estimate the value and gradient of the expected log joint. We use reparametrization tricks [Kingma and Welling, 2013] to get an unbiased estimate for the gradient of the expected log joint value. Apart from the difference in computing the expected log joint, all the other steps (optimization iterations, variational distribution initialization, etc.) stay the same as the variational inference part in vsbq.
C.5 MAP estimates via Bayesian Adaptive Direct Search (BADS)
@@
We provide the results for vsbq with cma-es and bads in Table S2, S3 and Figure S2, S3. From the tables and figures, we can see that the performance of vsbq with bads and cma-es is almost identical for three out of four benchmark problems (Two Moons, multivariate Rosenbrock-Gaussian, and Bayesian timing model). Instead, we found that vsbq performs less effectively with traces from bads in the multisensory causal inference model, compared to cma-es. Still, vsbq (bads) is comparable to the noiseless Laplace approximation (laplace) even in the presence of large amounts of log-density evaluation noise.
@@
Figure S2: Performance on Bayesian timing model with traces from different optimization algorithms. Median $\Delta$LML loss (left), MMTV (middle), and GsKL (right) as a function of the log-likelihood noise $\sigma_{\text{obs}}$ for the Bayesian Timing model. Shaded areas are 95% CI of the median and grey dash-dotted horizontal lines are the rule-of-thumb thresholds for good performance ($\Delta$LML=1, MMTV =0.2, GsKL =$1/8$). The performance of vsbq is virtually identical, regardless of the source of evaluations, whether cma-es or bads.
Figure S3: Performance on multisensory causal inference model with traces from different optimization algorithms. Median $\Delta$LML loss (left), MMTV (middle), and GsKL (right) as a function of the log-likelihood noise $\sigma_{\text{obs}}$ for the Bayesian Timing model. Shaded areas are 95% CI of the median and grey dash-dotted horizontal lines are the rule-of-thumb thresholds for good performance ($\Delta$LML=1, MMTV =0.2, GsKL =$1/8$). vsbq with traces from bads performs worse than vsbq with cma-es, but still comparable to the (noiseless) laplace.
C.6 Visualization of posteriors
We visualize posterior distributions as ‘corner plots’, i.e., a plot with 1D and all pairwise 2D marginals. For visualization of individual posteriors obtained by the algorithms, for all problems, we report example solutions obtained from a run with the same random seed (see Figure S5, S9, S13, and S17).
@@ C.6 Visualization of posteriors
(d) bbvi, Gaussian (diagonal)
(a) bbvi, Gaussian (full-rank)
(b) bbvi, MoG ($K=5$)
(c) bbvi, MoG ($K=50$)
Figure S5: Two Moons bimodal posterior visualization. The orange density contours and points in the sub-figures represent the posterior samples from vsbq, nnr, and laplace, while the black contours and points denote ground truth samples. vsbq and nnr perfectly capture the ground-truth bimodal posterior, while laplace is limited to one mode. Among the bbvi methods, the MoG (K=50) configuration performs best.
@@
(b) bbvi, Gaussian (diagonal)
(a) bbvi, Gaussian (full-rank)
(b) bbvi, MoG ($K=5$)
(a) bbvi, MoG ($K=50$)
Figure S9: Multivariate Rosenbrock-Gaussian posterior visualization. The orange density contours and points in the sub-figures represent the posterior samples from vsbq, nnr, and laplace, while the black contours and points denote ground truth samples. Both vsbq and nnr capture very well the complex shape of the distribution, while in this example laplace fails. Among the bbvi methods, the MoG (K=50) configuration performs best.
@@
(b) bbvi, Gaussian (diagonal)
(a) bbvi, Gaussian (full-rank)
(b) bbvi, MoG ($K=5$)
(a) bbvi, MoG ($K=50$)
Figure S13: Bayesian timing model posterior visualization. $\sigma_{\textrm{obs}}=3$ for vsbq and nnr, $\sigma_{\textrm{obs}}=0$ for laplace. The orange density contours and points in the sub-figures represent the posterior samples from vsbq, nnr, and laplace, while the black contours and points denote ground truth samples. Both vsbq and nnr capture the shape of the posterior in the presence of observation noise. laplace obtains a reasonable approximation, for the noiseless case. Among the bbvi methods, the Gaussian (diagonal), MoG (K=5), and MoG (K=50) configurations produce visually similar results and underperform compared to vsbq and nnr.
@@
(b) bbvi, Gaussian (diagonal)
(a) bbvi, Gaussian (full-rank)
(b) bbvi, MoG ($K=5$)
(a) bbvi, MoG ($K=50$)
Figure S17: Multisensory causal inference model posterior visualization. $\sigma_{\textrm{obs}}=3$ for vsbq and nnr, $\sigma_{\textrm{obs}}=0$ for laplace. The orange density contours and points in the sub-figures represent the posterior samples from vsbq, nnr, and laplace, while the black contours and points denote ground truth samples. $(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})=(\sigma_{\text{vis}}(c_{\text{low}}),\sigma_{\text{vis}}(c_{\text{med}}),\sigma_{\text{vis}}(c_{\text{high}}),\sigma_{\text{vest}},\lambda,\kappa)$. Both vsbq and nnr obtain a reasonable approximation of the complex posterior under noisy evaluations, with nnr yielding a more faithful approximation for this random seed. laplace fails to capture the posterior well, despite it being noiseless. Among the bbvi methods, the Gaussian (diagonal), MoG (K=5), and MoG (K=50) configurations produce visually similar results and underperform compared to vsbq and nnr.
C.7 Runtime analysis
In this section, we present an analysis of the runtime for our method and baselines. Such an analysis is important for a post-process inference method, which ideally should only take a relatively short time (e.g., a few minutes).333bbvi methods are not considered post-processing inference techniques and serve only as a reference baseline for posterior approximation accuracy. Their runtime primarily depends on the total number of target function evaluations and the grid search for tuning learning rates. For vsbq and nnr, we report the wall-clock runtime based on 5 independent runs with different training data sets (i.e., traces from the MAP optimizations). For laplace, the runtime is based on 5 runs for numerically estimating the Hessian matrix. We measure the runtime on both CPU and GPU platforms. [^4] It is worth noting that, in the case of GPU runtime measurement, only the sparse GP and neural network fitting were carried out with a GPU. Due to our current implementation, the stochastic variational inference optimization part for vsbq and nnr was always conducted using the CPU. Future efforts to port the stochastic variational inference procedure completely to GPU could further enhance the computational performance of both methods. Specifically, we utilize an AMD EPYC 7452 32-core Processor for CPU computations and an NVIDIA A100 for GPU computations.